OpenStudy (anonymous):

PLEASE HELP Evaluate the expression below when g = -6 -3g minus 2 minus(-4g + 10) or -3g - 2 -(-4g + 10)

OpenStudy (anonymous):

ugh

jimthompson5910 (jim_thompson5910):

-3g - 2 -(-4g + 10) -3(-6) - 2 -(-4(-6) + 10) 18 - 2 -(24 + 10) 18 - 2 -(34) 18 - 2 - 34 16 - 34 -18 So -3g - 2 -(-4g + 10) = -18 when g = -6

OpenStudy (anonymous):

thx jim i think that made more sense to me the other person knew the idea but i could tell that the way he solved it was not correct so i was confused

jimthompson5910 (jim_thompson5910):

if you have a graphing/scientific calculator (like a TI 83), then you can type the second line exactly as you see it and you'll be able to confirm the answer

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

9 squared is 81 and NOT 18

jimthompson5910 (jim_thompson5910):

and you square the 9 before you multiply

OpenStudy (anonymous):

k give me a second to try to re due that haha

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

hey i just found my TI-83 Plus silver edition and i was just curious if you know how to make the little two thing up above the other numbers ya know

jimthompson5910 (jim_thompson5910):

to find \[\Large 9^2\] type 9^2

jimthompson5910 (jim_thompson5910):

there is a ^ key on your calculator

OpenStudy (anonymous):

haha that was strangly obvious haha

jimthompson5910 (jim_thompson5910):

lol that's ok, if you've never used it before, then it's a bit strange at first and hard to figure out

OpenStudy (anonymous):

ok is the answer 180?

jimthompson5910 (jim_thompson5910):

it is, nice job

OpenStudy (anonymous):

awesome it feel really good to finally understand what the hell im doing lol and i think finding my calculator will be a huge help

jimthompson5910 (jim_thompson5910):

lol it definitely is a huge help and it can do a lot, so you'll be using it later down the road for sure

jimthompson5910 (jim_thompson5910):

and that's good you now know what's going on and such

OpenStudy (anonymous):

haha but i already have another question for you lol

jimthompson5910 (jim_thompson5910):

go for it

OpenStudy (anonymous):

i just tried to find the sum of this problem 4x^2+7x^2 but i got .498 and i dont think thats right the only confusing thing is i did it on the calculator and i dont know how i put it in wrong on the calculator

jimthompson5910 (jim_thompson5910):

4x^2 and 7x^2 are like terms because they have the same variable with the same exponent

jimthompson5910 (jim_thompson5910):

so you can add the coefficients: 4+7 = 11

jimthompson5910 (jim_thompson5910):

this means 4x^2+7x^2 = 11x^2

OpenStudy (anonymous):

so when i am looking at an equation how do i just know that they are like terms

jimthompson5910 (jim_thompson5910):

Notice they both have an x

jimthompson5910 (jim_thompson5910):

And the exponent over the x is 2

jimthompson5910 (jim_thompson5910):

if you had 4x^2 and 7x, then they would NOT be like terms

OpenStudy (anonymous):

oh obviously still blonde

jimthompson5910 (jim_thompson5910):

so it's making sense now?

OpenStudy (anonymous):

is it 11 x^2

jimthompson5910 (jim_thompson5910):

it is

OpenStudy (anonymous):

ok so am i able to figure that out on the calculator

jimthompson5910 (jim_thompson5910):

not with a TI 83, but you can add the coefficients

jimthompson5910 (jim_thompson5910):

you can add 4+7 on the calculator, but I'm betting you can probably do that in your head

OpenStudy (anonymous):

haha ok :)